SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Aug 16, 2005)

Attendees: Yaw Ming Chia, Al Eisner, Fang Fang, Selina Li, Brian Lindquist, Dan Walker.

Operations

  • The monthly TRG status report was given last Wednesday (Aug.10). Click here to see the talk.
  • We have been testing the new GLT firmware at IR2 for the past week. See the GLT section below.
  • Online release 13.4.0 based on offline release 18.2.1 was installed at IR2 today. This release included the fix for the L3 node trim due to L3 tracking.

EMT

  • Dan is making progress in saving and retrieving the prescale from his own database. The plan is to test the ROM code at the test stand sometime on Thursday or Friday of this week.

DCT/DCZ

  • We got paged on Monday owl shift about dataflow errors for damage 4000 from DCZ which prevented ORC from configuring for data-taking. The master crate was rebooted due to PEP fiducial timing problem. Normally if it's a reboot, we don't need to power cycle the DCZ crates, but this time power-cycling the DCZ crates solved the problem with ORC not being able to configure due to interruption of timing to the DCZ clock.
  • There are some new warning messages found in the TSF-Y ROM logfile which is located here. See HN for more detail. In this HN, there is also discussions about problem with the "broken" power switch for the TSF-Y crate.
  • Fang will be updating the online DCT/DCZ ROM lib to the new online release after our operation meeting.
  • Ming has been gathering data/information for Marc to understand the problem with TSF crate power-cycling (at the test stand).
  • This is going to be Ming's last time at the trigger operation meeting as he will be leaving us and going back to UK this weekend. There is no one currently taking over his duties for the TSF.

GLT

  • One physics run (57087) was taken last Friday with a new version of the GLT firmware in. There were EMT->GLT and DCT->GLT transmission errors, as well as GLT-FC word corruptions. This version had about 1% event problem that could be due to the DAQ buffer tracking occassionally slips or something fundamental in the DAQ FPGA firmware that actually cannot deliver the DAQ data fast enough under some circumstances.
  • Su Dong left us 2 versions of the firmware for us to test this week to target these two possibilities mentioned above.

L3

  • There was one L3 node trim on Sunday, and the node recovered itself after the next run.
  • There should be no more L3 node trim due to L3 tracking since the code was restored in this online release today (Aug.16, Tuesday).
  • Prafulla is out of the country right now. Currently, no one is taking the L3 pager. Arik is possibly taking over Prafulla's L3 tasks while Prafulla is away.

Data Quality and Monitoring

  • Nothing to report.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.